Hello. I have a question about triangles. My problem is:
The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the length of the base. If the perimeter is 29cm, find the length of one of the equal legs.
I have tried everyother website I could find online and they all seem to lean more towards right triangles. Thank you for your help.

Answer by stanbon(75887)

The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the length of the base. If the perimeter is 29cm, find the length of one of the equal legs.
Perimeter = leg + leg + base
29=base + (4base-8)+(4base-8)
29=b+8b-16
9b=45
b=5 cm (This is the length of the base)
Cheers,
Stan H.